OpenStudy (anonymous):

What is the equation of the line in standard form that passes through the point (–2, –3) and is parallel to the line y = x + 9?

OpenStudy (aroub):

parallel=same slope y=mx+b , m= 1 , ( -2,-2) -2=1(-2)+b -2+2=b 0=b y=x

OpenStudy (anonymous):

x – y = 1 x + y = 1 x – y = –5 x + y = 5

OpenStudy (anonymous):

y=x-1............subract x from both sides -x+y=-1............multiply by -1 to the equation to change the signs x-y=1

OpenStudy (anonymous):

yes a. x-y=1